OpenStudy (joshoyen):

Complete the square then rewrite as a perfect square. HELP PLS 1. z^2 - 19z + ? = 0

OpenStudy (anonymous):

takes two steps what is half of 19 (answer as a fraction)

OpenStudy (joshoyen):

oh wait i got that part i just need help rewriting it as a perfect square

OpenStudy (joshoyen):

it is z^2-19z+90.25=0

OpenStudy (anonymous):

actually question is unclear but maybe they just want \[\left(z-\frac{19}{2}\right)^2=0\]

OpenStudy (joshoyen):

wait no you got it right, my answer sheet says 361/4; (z-19/2)^2

OpenStudy (anonymous):

ok

OpenStudy (joshoyen):

how did you get that though, i dont know the steps into rewriting it as ap erfect square

OpenStudy (anonymous):

half of \(19\) is \(\frac{19}{2}\) and \(\left(\frac{19}{2}\right)^2=\frac{361}{4}\)

OpenStudy (anonymous):

don't muck around with decimals

OpenStudy (anonymous):

Actually it tends to be confusing, what really is the value of that question mark? Is it another variable or it's just the same as z or it has a numerical value?

OpenStudy (joshoyen):

wait so what would be the rewritten perfect square form for a^2+14a+49=0?

OpenStudy (anonymous):

half of 14 is 7 and \(7^2=19\) so \[a^2+14a+49=(a+7)^2\]

OpenStudy (anonymous):

whoops meant to say \(7^2=49\) but you get the idea

OpenStudy (joshoyen):

alright thanks a lot!

OpenStudy (joshoyen):

Wait can I get one more problem? What would p^2 + 3p + ? be?